OpenStudy (anonymous):

what is sqrt{9} + sqrt{98} + sqrt{81} + sqrt{32}?

OpenStudy (anonymous):

\[\sqrt{9} + \sqrt{98} +\sqrt{81} +\sqrt{32}\]

OpenStudy (anonymous):

27.556

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

sqrt9 = 3 sqrt98 = sqrt(49*2) = 7sqrt2 sqrt81 = 9 sqrt32 = sqrt(16*2) = 4sqrt2 so adding them all up: 3 + 9 + 7sqrt2 + 4sqrt2 = 12 + 11sqrt2